# Membership Inference Attacks and Privacy in Topic Modeling

### v1

@@ 1 Introduction
Deep learning models’ propensity to memorize training data presents many privacy concerns. Notably, large language models are particularly susceptible to privacy attacks that exploit memorization [6, 5]. In this paper, we aim to investigate whether simpler probabilistic models, such as topic models, raise similar concerns. Our specific focus lies on examining probabilistic topic models like Latent Dirichlet Allocation (LDA) [2].
@@
Topic modeling is an unsupervised machine learning (ML) method that aims to identify underlying themes or topics within a corpus. Despite the recent successes of large language models (LLMs), Probabilistic topic models are still widely used due to their interpretability and straightforward implementation for text analysis. Furthermore, topic models continue to be developed for a variety of applications and downstream tasks like document classification, summarization, or generation [3, 32].
@@
Researchers apply topic models across a variety of domains where privacy concerns arise. For example, many studies in the medical domain use topic models to analyze data sets with sensitive attributes [21]. Additionally, topic models are widely used for various government or national defense applications. For instance, researchers applied LDA to better understand Twitter activity aimed at discrediting NATO during the Trident Juncture Exercises in 2018 [29]. To ensure ethical and responsible use of ML, it is crucial to consider the privacy implications associated with topic modeling.
@@
Probabilistic topic models like LDA serve as a basic Bayesian generative model for text. Neural topic models and language models can learn complex generative processes for text based on observed patterns in the training data. However, LDA requires the generative process for documents to be explicitly defined a priori with far fewer parameters and a Bag-of-Words (BoW) text representation. Investigating LDA’s susceptibility to privacy attacks provides broader context for the privacy vulnerabilities researched in contemporary ML models. Additionally, our work informs practitioners who may opt to use simpler topic models under an unjustified impression that they do not share the same vulnerabilities of LLMs.
@@
To explore the privacy in topic modeling, we conduct membership inference attacks (MIAs) which infer whether or not a specific document was used to train LDA. We propose an attack based on an LDA-specific query statistic designed to exploit memorization. This query statistic is integrated into the Likelihood Ratio Attack (LiRA) framework introduced by Carlini et al. [4]. We show that our attack can confidently infer the membership of documents included the training data of LDA which indicates that the privacy risks in generative modeling are not restricted to large neural models.
@@
To mitigate such vulnerabilities, we explore differential privacy (DP) for topic modeling. DP acts a direct defense against MIAs by providing a statistical guarantee that the output of some data analysis is indistinguishable regardless of any one user’s inclusion in the analysis. While several DP topic modeling algorithms in the literature attempt to protect individual privacy, previous works largely disregard the privacy of the model’s accompanying vocabulary set [33, 19, 18, 9, 22, 31]. Instead, they consider the vocabulary set as given or public information. However, in practice, the vocabulary set is derived from the training data and can leak sensitive information. We propose an algorithm for DP topic modeling provides privacy in both the vocabulary selection and learning method. By incorporating DP vocabulary selection into the private topic modeling workflow, our algorithm enhances privacy guarantees and bolsters defenses against the LiRA, while having minimal impact on practical utility.
@@ 2 Background and Notation
We let $f_{\cal{G}}(D)$ denote the process of learning the model parameterized by latent variables in $\mathcal{G}$ on a corpus $D\in[0,\infty]^{M\times V}$ with $M$ documents drawn from the underlying data distribution $\mathbb{D}$. The function $f_{\mathcal{G}}$ incorporates the learning algorithm to estimate $\mathcal{G}$’s latent variables and returns the learned topic model $\Phi$. Our definition assumes that the primary objective of topic modeling is to estimate $\Phi$. While $\mathcal{G}$ is also associated with other latent variables, $\Phi$ best achieves the goals associated with topic modeling by summarizing the relevant themes in the corpus.
@@ 2.1 Latent Dirichlet Allocation
Blei et al. [2] define LDA where each document is assumed to contain a mixture of topics. LDA assumes that each document $d\in D$ is represented by a $k$-dimensional document-topic distribution $\theta\sim\text{Dirichlet}(\alpha)$, and the entire corpus is represented by $k$ topic-word distributions $\phi\sim\text{Dirichlet}(\beta)$. Furthermore, it models each word in a document as generated by sampling a topic from the document’s topic distribution and then sampling a word from the chosen topic’s distribution over words.
The process of estimating $\Phi$ presents a Bayesian inference problem typically solved using collapsed Gibbs sampling or variational inference [14, 15]. Each entry $\Phi_{z,w}$ represents the probability of drawing word $w$ from topic $z$. The entire topic model $\Phi$ represents the likelihood of each word appearing in each topic, which can be used to estimate the document-topic distribution $\theta$ for any given document.
@@ 2.2 Memorization and Privacy
Understanding memorization in machine learning is critical for trustworthy deployment. Neural model’s large parameter space and learning method introduce training data memorization which increases with model size or observation duplication [5, 12]. Memorization introduces vulnerabilities to attacks on privacy, such as membership inference, attribute inference or data extraction attacks [25, 27, 6].
Schofield et al. [23] note that document duplication in topic models concentrates the duplicated document’s topic distribution $\theta$ over less topics, and briefly refer to this phenomenon as memorization. Their findings are consistent with studies on memorization and text duplication in large language models [5]. In this work, we investigate the memorization and privacy of topic models using membership inference attacks (MIA).
@@ 2.3 Membership Inference Attacks
In an MIA the adversary learns if a specific observation appeared in the model’s training data [16, 25]. These attacks serve as the foundation for other attacks like data extraction, and could violate privacy alone if the adversary learns than an individual contributed to a data set with a sensitive global attribute (i.e. a data set containing only patients with disease X).
Prior to attacking ML models, researchers explored exploiting aggregate statistics released in genome-wide association studies (GWAS) [16]. Shokri et al. [25] introduced one of the first MIAs on ML models. While most MIAs exploit deep learning models, few studies investigate topic models. Huang et al. [18] propose three simple MIAs to evaluate their privatized LDA learning algorithm. However, they evaluate their attacks using precision and recall and fail to show that their attack confidently identify members of the training data.
To evaluate an MIA’s ability to confidently infer membership, we examine the attack’s true positive rate (TPR) at low false positive rates (FPR) [4]. Receiver operator characteristic (ROC) and area under the curve (AUC) analysis is useful, but we must be wary of the axes. To accurately capture the TPR at the low FPRs of interest, we visualize ROC curves using log-scaled axes.
@@ 3.1 Threat Model
Under the typical query model, we could consider queries as a request to predict a document-topic distribution $\hat{\theta}$ given a document. However, $\hat{\theta}_{d}$ would hold little meaning without access to $\Phi$. Furthermore, a clever adversary could query many specially selected documents to reconstruct $\Phi$. Therefore, we remove this level of abstraction and assume that the adversary is given direct access to $\Phi$ as the output of our topic model.
@@ 3.2 Attack Framework
In the LiRA framework of Carlini et al. [4], the adversary performs hypothesis testing to determine whether or not a target document $d$ was in the training data. Specifically, we consider $\textbf{T}_{in}(d)=\{\Phi\leftarrow f_{\mathcal{G}}(D\cup\{d\})|D\in\mathbb{D}\}$ and $\textbf{T}_{out}(d)=\{\Phi\leftarrow f_{\mathcal{G}}(D\backslash\{d\})|D\in\mathbb{D}\}$ where $\textbf{T}_{in}(d)$ is the distribution of $\Phi$ learned with $d$, and $\textbf{T}_{out}(d)$ is the distribution of $\Phi$ learned without $d$. Given an observed $\Phi_{obs}$, would like to construct the likelihood ratio test statistic $\Lambda$ as follows:
@@
where $p(\Phi_{obs}|\textbf{T}_{x}(d))$ is the probability density function of $\Phi_{obs}$ under $\textbf{T}_{x}(d)$ [4].
However, the ration in Equation 1 is infeasible to calculate directly because all $\Phi$ represent a set of $k$ $V$-dimensional distributions. Instead, the adversary applies a carefully chosen query statistic $\zeta:(\Phi,d)\rightarrow\mathbb{R}$ on all $\Phi$ to reduce the problem to 1-dimension. The test proceeds by calculating the probability density of $\zeta(\Phi_{obs},d)$ under estimated normal distributions from $\tilde{\textbf{T}}_{in}(d)=\{\zeta(\Phi,d)\ |\ \Phi\leftarrow f_{\mathcal{G}}(D\cup\{d\}),D\leftarrow\mathbb{D}\}$ and $\tilde{\textbf{T}}_{out}(d)=\{\zeta(\Phi,d)\ |\ \Phi\leftarrow f_{\mathcal{G}}(D\backslash\{d\}),D\leftarrow\mathbb{D}\}$:
@@
In the online LiRA, the adversary must train $N$ shadow topic models with and without $d$ to estimate $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$. Appendix A contains the full online LiRA algorithm. To ease the computational burden of training $N$ shadow topic models for each target document, Carlini et al. propose an offline variant of the LiRA. In the offline LiRA, the adversary can evaluate any $d$ after learning a collection of $N$ shadow topic models once by comparing $\zeta(\Phi_{obs},d)$ to a normal estimated from $\tilde{\textbf{T}}_{out}(d)$ with a one-sided hypothesis test. Appendix B contains the full offline LiRA algorithm.
Designing $\zeta$ is crucial to the attack’s success. In supervised learning scenarios, the adversary can directly query the model and use the loss of an observation to inform the statistic [4]. However, because topic models are unsupervised learning tasks, the adversary encounters the challenge of identifying an informative statistic that can be computed efficiently using $\Phi$. Hence, $\zeta$ must be careful tailored to topic models for the attack to be effective.
@@ 3.3 Designing an Effective Query Statistic $\zeta$
Effective model query statistics for the LiRA framework must satisfy the following criteria:
  1. 1.
  2. 2.
Ideally, the statistic should also be related to how the model may memorize the training data to provide an intuitive interpretation for attack performance.
First, we consider statistics from the simple MIAs on LDA presented by Huang et al. [18]. They propose three statistics derived from the target document’s estimated topic distribution $\hat{\theta}=\hat{\theta}(\Phi_{obs},d)$: the entropy of $\hat{\theta}$, the standard deviation of $\hat{\theta}$, and the maximum value in $\hat{\theta}$. The intuition behind their chosen statistics comes from the observation that a document’s topic distribution tends to concentrate in a few topics when included (or duplicated) in the training data [23]. However, Huang et al. apply global thresholds to these statistics instead of using them as query statistics to derive the LiRA test statistic as in Equation 2.
@@
where we optimzie over $\theta\in[0,1]^{k}$ s.t. $\sum_{z}\theta_{z}=1$. In practice, we use SciPy’s out-of-the-box optimization methods to estimate $\zeta(\Phi,d)$ [30].
The intuition behind our statistic is that a document is more likely generated by the target model when included the training data. Therefore, using $\zeta$ allows us to reason about memorization in topic models. We verify our criteria for the proposed statistic in Appendix D. Additionally, Appendix D contains an evaluation of other candidate statistics where we show that our statistic significantly outperforms other statistics based on [18]
The LiRA with statistic $\zeta$ directly exploits memorization and per-example hardness. The attack accounts for the natural differences in $\zeta$ on various documents by estimating the likelihood ratio under distributions based on $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$. This enables the LiRA to outperform simple attacks like in [18] that use global thresholds on the other query statistics based on $\hat{\theta}$. Therefore, we propose the LiRA with statistic $\zeta$ as a stronger alternative to attack topic models.
@@ 3.4 Memorization and Per-Example Hardness
To verify aspects of memorization and per-example hardness for probabilistic topic models, we estimate and visualize $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$ in an experiment similar to [4] and [12]. Consistent with their findings, we show that outlying and hard-to-fit observations tend to have a large effect on the learned model when included in the training set for LDA. The histograms in Figure 1 display the estimated distributions for various types of documents.
@@
The simple fact that model is more likely to generate specific documents after inclusion in the training data hints toward memorization. Because the generative process for documents in LDA is extremely simple compared to language models, they do not learn verbatim sequences of words. Instead, topic model’s ability to “memorize" the training data is due to learning based on word co-occurrences.
Long document’s are inherently harder to fit than shorter documents. Because the log-likelihood for a document $d$ is the sum of independent log-probabilities for generating each word in $d$, the magnitude of $\zeta(\Phi,d)$ is naturally higher for longer documents. Additionally, the model may struggle to capture coherent topic structures for longer documents because they tend to contain words from a wider range of topics.

### v2

@@ 1 Introduction
Deep learning models’ propensity to memorize training data presents many privacy concerns. Notably, large language models are particularly susceptible to privacy attacks that exploit memorization (Carlini et al., 2021; 2023). In this paper, we aim to investigate whether simpler probabilistic models, such as topic models, raise similar concerns. Our specific focus lies on examining probabilistic topic models like Latent Dirichlet Allocation (LDA) (Blei et al., 2003).
@@
Topic modeling is an unsupervised machine learning (ML) method that aims to identify underlying themes or topics within a corpus. Despite the recent successes of large language models (LLMs), Probabilistic topic models are still widely used due to their interpretability and straightforward implementation for text analysis. Furthermore, topic models continue to be developed for a variety of applications and downstream tasks like document classification, summarization, or generation (Boyd-Graber et al., 2017; Wang et al., 2019).
@@
Researchers apply topic models across a variety of domains where privacy concerns arise. For example, many studies in the medical domain use topic models to analyze datasets with sensitive attributes (Mustakim et al., 2021). Additionally, topic models are widely used for various government or national defense applications. For instance, researchers applied LDA to better understand Twitter activity aimed at discrediting NATO during the Trident Juncture Exercises in 2018 (Uyheng et al., 2020). To ensure ethical and responsible use of ML, it is crucial to consider the privacy implications associated with topic modeling.
@@
Probabilistic topic models like LDA serve as a basic Bayesian generative model for text. Neural topic models and language models can learn complex generative processes for text based on observed patterns in the training data. However, LDA requires the generative process for documents to be explicitly defined a priori with far fewer parameters and a Bag-of-Words (BoW) text representation. Investigating LDA’s susceptibility to privacy attacks provides broader context for the privacy vulnerabilities researched in more complex ML models. Additionally, our work informs practitioners who may opt to use topic models under the false impression that they do not share the same vulnerabilities of LLMs.
@@
To explore the privacy in topic modeling, we conduct membership inference attacks (MIAs) which infer whether or not a specific document was used to train LDA. We propose an attack based on an LDA-specific query statistic designed to exploit memorization. This query statistic is integrated into the Likelihood Ratio Attack (LiRA) framework introduced by Carlini et al. (2022). We show that our attack can confidently infer the membership of documents included in the training data of LDA which indicates that the privacy risks in generative modeling are not restricted to large neural models.
@@
To mitigate such vulnerabilities, we explore differential privacy (DP) for topic modeling. DP acts a direct defense against MIAs by providing a statistical guarantee that the output of some data analysis is indistinguishable regardless of any one user’s inclusion in the analysis. While several DP topic modeling algorithms in the literature attempt to protect individual privacy, previous works largely disregard the privacy of the model’s accompanying vocabulary set (Zhao et al., 2021; Huang & Chen, 2021; Huang et al., 2022; Decarolis et al., 2020; Park et al., 2016; Wang et al., 2022). Instead, they consider the vocabulary set as given or public information. However, in practice, the vocabulary set is derived from the training data and can leak sensitive information. We propose an algorithm for DP topic modeling that provides privacy in both the vocabulary selection and learning method. By incorporating DP vocabulary selection into the private topic modeling workflow, our algorithm enhances privacy guarantees and bolsters defenses against the LiRA, while having minimal impact on practical utility.
@@ 2 Background and Notation
We let $f_{\cal{G}}(D)$ denote the process of learning a model parameterized by latent variables in $\mathcal{G}$ on a corpus $D$ with $M$ documents drawn from the underlying data distribution $\mathbb{D}$. The function $f_{\mathcal{G}}$ incorporates the learning algorithm to estimate $\mathcal{G}$’s latent variables and returns the learned topic model $\Phi$. Our definition assumes that the primary objective of topic modeling is to estimate $\Phi$. While $\mathcal{G}$ is also associated with other latent variables that may define the document generation process, $\Phi$ best achieves the goals associated with topic modeling by summarizing the relevant themes in the corpus.
@@ 2.1 Latent Dirichlet Allocation
Blei et al. (2003) define LDA where each document is assumed to contain a mixture of topics. LDA assumes that each document $d\in D$ is represented by a $k$-dimensional document-topic distribution $\theta\sim\text{Dirichlet}(\alpha)$, and the entire corpus is represented by $k$ $V$-dimensional topic-word distributions $\phi\sim\text{Dirichlet}(\beta)$. Furthermore, it follows a simple bag-of-words approach where each word in a document is generated by first sampling a topic from the document’s topic distribution and then sampling a word from the chosen topic’s distribution over words.
The process of estimating $\Phi$ presents a Bayesian inference problem typically solved using collapsed Gibbs sampling or variational inference (Griffiths & Steyvers, 2004; Hoffman et al., 2010). Each entry $\Phi_{z,w}$ represents the probability of drawing word $w$ from topic $z$. The entire topic model $\Phi$ represents the likelihood of each word appearing in each topic, which can be used to estimate the document-topic distribution $\theta$ for any given document.
@@ 2.2 Memorization and Privacy
Understanding memorization in machine learning is critical for trustworthy deployment. Neural model’s large parameter space and learning method introduce training data memorization which increases with model size or observation duplication (Carlini et al., 2023; Feldman & Zhang, 2020). Memorization introduces vulnerabilities to attacks on privacy, such as membership inference, attribute inference or data extraction attacks (Shokri et al., 2017; Song & Raghunathan, 2020; Carlini et al., 2021).
Schofield et al. (2017) note that document duplication in topic models concentrates the duplicated document’s topic distribution $\theta$ over less topics, and briefly refer to this phenomenon as memorization. Their findings are consistent with studies on memorization and text duplication in large language models (Carlini et al., 2023). In this work, we investigate the memorization and privacy of topic models using membership inference attacks (MIA).
@@ 2.3 Membership Inference Attacks
In an MIA the adversary learns if a specific observation appeared in the model’s training data (Homer et al., 2008; Shokri et al., 2017). These attacks serve as the foundation for other attacks like data extraction, and could violate privacy alone if the adversary learns that an individual contributed to a dataset with a sensitive global attribute (i.e. a dataset containing only patients with disease X).
Prior to attacking ML models, researchers explored exploiting aggregate statistics released in genome-wide association studies (GWAS) (Homer et al., 2008). Shokri et al. (2017) introduced one of the first MIAs on ML models. While most MIAs exploit deep learning models, few studies investigate topic models. Huang et al. (2022) propose three simple MIAs to evaluate their privatized LDA learning algorithm. However, they evaluate their attacks using precision and recall and fail to show that their attack confidently identifies members of the training data.
To evaluate an MIA’s ability to confidently infer membership, we examine the attack’s true positive rate (TPR) at low false positive rates (FPR) (Carlini et al., 2022). Receiver operator characteristic (ROC) and area under the curve (AUC) analysis may be useful, but we must be wary of the axes. To accurately capture the TPR at the low FPRs of interest, we visualize ROC curves using log-scaled axes.
@@ 3.1 Threat Model
Under the typical query model, we could consider queries as a request to predict a document-topic distribution $\hat{\theta}$ given a document. However, $\hat{\theta}$ would hold little meaning without access to $\Phi$. Furthermore, a clever adversary could query many specially selected documents to reconstruct $\Phi$. Therefore, we remove this level of abstraction and assume that the adversary is given direct access to $\Phi$ as the output of our topic model.
@@ 3.2 Attack Framework
In the LiRA framework of Carlini et al. (2022), the adversary performs hypothesis testing to determine whether or not a target document $d$ was in the training data. Specifically, we consider $\textbf{T}_{in}(d)=\{\Phi\leftarrow f_{\mathcal{G}}(D\cup\{d\})|D\in\mathbb{D}\}$ and $\textbf{T}_{out}(d)=\{\Phi\leftarrow f_{\mathcal{G}}(D\backslash\{d\})|D\in\mathbb{D}\}$ where $\textbf{T}_{in}(d)$ is the distribution of $\Phi$ learned with $d$, and $\textbf{T}_{out}(d)$ is the distribution of $\Phi$ learned without $d$. Given an observed $\Phi_{obs}$, we would like to construct the likelihood ratio test statistic $\Lambda$ as follows:
@@
where $p(\Phi_{obs}|\textbf{T}_{x}(d))$ is the probability density function of $\Phi_{obs}$ under $\textbf{T}_{x}(d)$ (Carlini et al., 2022).
However, the ratio in Equation 1 is intractable because it requires integrating over all $\Phi$ with all possible $D$ with and without $d$. Instead, the adversary applies a carefully chosen query statistic $\zeta:(\Phi,d)\rightarrow\mathbb{R}$ on all $\Phi$ to reduce the problem to 1-dimension. The test can proceed by calculating the probability density of $\zeta(\Phi_{obs},d)$ under estimated normal distributions from $\tilde{\textbf{T}}_{in}(d)=\{\zeta(\Phi,d)\ |\ \Phi\leftarrow f_{\mathcal{G}}(D\cup\{d\}),D\leftarrow\mathbb{D}\}$ and $\tilde{\textbf{T}}_{out}(d)=\{\zeta(\Phi,d)\ |\ \Phi\leftarrow f_{\mathcal{G}}(D\backslash\{d\}),D\leftarrow\mathbb{D}\}$:
@@
In the online LiRA, the adversary must train $N$ shadow topic models with and without $d$ to estimate $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$. Appendix A contains the full online LiRA algorithm. To ease the computational burden of training $N$ shadow topic models for each target document, Carlini et al. (2022) propose an offline variant of the LiRA. In the offline LiRA, the adversary can evaluate any $d$ after learning a collection of $N$ shadow topic models once by comparing $\zeta(\Phi_{obs},d)$ to a normal estimated from $\tilde{\textbf{T}}_{out}(d)$ with a one-sided hypothesis test. Appendix B contains the full offline LiRA algorithm.
Designing $\zeta$ is crucial to the attack’s success. In supervised learning scenarios, the adversary can directly query the model and use the loss of an observation to inform the statistic (Carlini et al., 2022). However, because topic models present an unsupervised learning tasks, the adversary encounters the challenge of identifying an informative statistic that can be computed efficiently using $\Phi$. Hence, $\zeta$ must be carefully tailored to topic models for the attack to be effective.
@@ 3.3 Designing an Effective Query Statistic $\zeta$
Effective model query statistics for the LiRA framework must satisfy the following criteria: the statistic should increase for $d$ when $d$ is included in the training data ($\tilde{\textbf{T}}_{in}(d)>\tilde{\textbf{T}}_{out}(d)$) to enable the offline LiRA, and the estimated distributions $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$ are approximately normal to allow for parametric modeling. Ideally, the statistic should be related to how the model may memorize the training data to provide an intuitive interpretation for attack performance.
First, we consider statistics from the simple MIAs on LDA presented by Huang et al. (2022). They propose three statistics derived from the target document’s estimated topic distribution $\hat{\theta}$: the entropy of $\hat{\theta}$, the standard deviation of $\hat{\theta}$, and the maximum value in $\hat{\theta}$. The intuition behind their chosen statistics comes from the observation that a document’s topic distribution tends to concentrate in a few topics when included (or duplicated) in the training data (Schofield et al., 2017). However, Huang et al. (2022) apply global thresholds to these statistics instead of using them as query statistics to derive the LiRA test statistic as in Equation 2.
@@
where we optimzie over $\theta\in[0,1]^{k}$ s.t. $\sum_{z}\theta_{z}=1$. In practice, we use SciPy’s out-of-the-box optimization methods to estimate $\zeta(\Phi,d)$ (Virtanen et al., 2020).
Compared to the statistics proposed by Huang et al. (2022), our statistic more directly exploits LDAs generative processes and is based on the observation that a document’s likelihood under the target model increases when included in the training data. Therefore, using $\zeta$ allows us to better reason about memorization in topic models. We verify our criteria for the proposed statistic in Appendix C. Additionally, Appendix C contains an evaluation of other candidate statistics where we show that our statistic significantly outperforms other statistics based on the attacks in Huang et al. (2022).
The LiRA with statistic $\zeta$ directly exploits memorization and per-example hardness. The attack accounts for the natural differences in $\zeta$ on various documents by estimating the likelihood ratio under distributions based on $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$. This enables the LiRA to outperform simple attacks like in Huang et al. (2022) that use global thresholds on the other query statistics based on $\hat{\theta}$. Therefore, we propose the LiRA with statistic $\zeta$ as a stronger alternative to attack topic models.
@@ 3.4 Memorization and Per-Example Hardness
To verify aspects of memorization and per-example hardness for probabilistic topic models, we estimate and visualize $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$ in an experiment similar to Carlini et al. (2022) and Feldman & Zhang (2020). Consistent with their findings, we show that outlying and hard-to-fit observations tend to have a large effect on the learned model when included in the training set for LDA. The histograms in Figure 1 display the estimated distributions for various types of documents.
@@
The simple fact that the model is more likely to generate specific documents after inclusion in the training data hints toward memorization. Because the generative process for documents in LDA is extremely simple compared to neural language models, they do not learn verbatim sequences of words. Instead, topic model’s ability to “memorize" the training data is due to learning based on word co-occurrences.
Long documents are inherently harder to fit than shorter documents. Because the log-likelihood for a document $d$ is the sum of independent log-probabilities for generating each word in $d$, the magnitude of $\zeta(\Phi,d)$ is naturally higher for longer documents. Additionally, the model may struggle to capture coherent topic structures for longer documents because they tend to contain words that span a wider range of topics.